OpenStudy (anonymous):

cos(45 °-30 °) I know how to set this up using sum and Difference Formula. But multiplying radical fractions make 0 sense to me. can someone explain how the answer is (1+sqrt(3))/(2 sqrt(2))

OpenStudy (anonymous):

\[\cos(A-B)=\cos Acos B+sinAsinB\]

OpenStudy (anonymous):

put A=45 and B=30

OpenStudy (anonymous):

\[(\sqrt{2}/\sqrt{2})(\sqrt{3}/2)+(\sqrt{2}/2)(1/2)= (\sqrt{2+3}/2) + (\sqrt{2}/4)\] Is this the correct process?

OpenStudy (phi):

you mean sqrt(2)/2 not sqrt(2)/sqrt(2) when you multiply fractions, you multiply top*top and bottom * bottom

OpenStudy (anonymous):

sorry for the incorrect input of my answers.... Im new to this

OpenStudy (anonymous):

\[(\sqrt{2}+\sqrt{3}/2) + (\sqrt{2}/4)\]

OpenStudy (phi):

If you use the equation editor, it's easier to put in the correct equation. but you start with \[ \frac{\sqrt{2}}{2} \cdot \frac{\sqrt{3}}{2} +\frac{\sqrt{2}}{2} \cdot \frac{1}{2} \]

OpenStudy (phi):

you did the right-hand side fraction multiply correctly: sqrt(2)*1 is sqrt(2) for the top and 2*2= 4 for the bottom try again for the left fractions

OpenStudy (phi):

btw, you can write either \[ \sqrt{2}\cdot\sqrt{3} \] or \[ \sqrt{2\cdot3}= \sqrt{6} \]

OpenStudy (phi):

so the answer is \[ \frac{\sqrt{2}}{2} \cdot \frac{\sqrt{3}}{2} +\frac{\sqrt{2}}{2} \cdot \frac{1}{2} = \frac{\sqrt{2}\cdot \sqrt{3} + \sqrt{2}}{4} \] you could "factor out" a sqrt(2) from each term in the top and write the expression as \[ \frac{\sqrt{2}(\sqrt{3} + 1)}{4} \] Though it is generally not done to put a square root in the denominator , you could multiply top and bottom by sqrt(2) \[ \frac{\sqrt{2}(\sqrt{3} + 1)}{4} \cdot \frac{ \sqrt{2}}{\sqrt{2}}\]

OpenStudy (phi):

can you finish ?

OpenStudy (anonymous):

no Im stuck! haha. I know the answer is : \[(1+\sqrt{3})/(2\sqrt{2})\] I dont understand how to get this answer

OpenStudy (phi):

for this \[ \frac{\sqrt{2}(\sqrt{3} + 1)}{4} \cdot \frac{ \sqrt{2}}{\sqrt{2}} \] you multiply top times top and bottom times bottom we can change the order of multiplication so we can write it as \[ \frac{\sqrt{2}\sqrt{2}(\sqrt{3} + 1)}{4\sqrt{2}} \] what is sqrt(2)*sqrt(2) ?

OpenStudy (phi):

you could use a calculator, but you should learn that sqrt(something)*sqrt(something)= something where the something inside the square root is the same thing 1?? is your calculator broken ?

OpenStudy (anonymous):

\[\sqrt{2}\times \sqrt{2}\]=1.682

OpenStudy (anonymous):

sorry I'm having the hardest time with this and it should be easy! thank you for your help

OpenStudy (phi):

\[ \sqrt{2}\times \sqrt{2} \] out of curiosity how did you get 1.682 ?

OpenStudy (phi):

ok you did \[ \sqrt{2 \sqrt{2}} \] which is not what you want. you could type this into the google search window sqrt(2)*sqrt(2)=

OpenStudy (anonymous):

ok it would be 2.

OpenStudy (phi):

yes, now what do we have for your problem after replacing sqrt(2)*sqrt(2) with 2

OpenStudy (anonymous):

\[\frac{ 2(\sqrt{3}+1) }{ 4\sqrt{2} }\]

OpenStudy (phi):

do you know you can simplify by dividing top and bottom by 2 (or multiply top and bottom by 1/2) what do you get ?

OpenStudy (phi):

in other words 2/4 is the same as 1/2

OpenStudy (anonymous):

\[ \frac{ \sqrt{3} +1}{ 2\sqrt{2} }\]

OpenStudy (anonymous):

after dividing by 2 thats what I got

OpenStudy (phi):

does that look familiar ?

OpenStudy (anonymous):

that's the final answer. I just don't feel comfortable doing these problems yet. Hopefully it will click before my test ! Thank you for your help!

OpenStudy (phi):

I would watch the Khan videos on radicals (and fractions if you need help there) There are only a few "rules" to learn, and then some practice so it sinks in.

OpenStudy (anonymous):

ok, im watching it now! forgot about this website, very helpful